# Impurity flat band states in the diamond chain

### v1

@@ IV.3.1 Topological characterization
This behavior is remarkably similar to that observed in chiral-symmetric non-Hermitian 1D models where, due to the fact that each exceptional point (EP) [50] carries a topological charge of ±$\pm 1/2$, the absolute value of the winding number, $|W|$, corresponds to half the number of times each EP is encircled as the momentum $k$ traverses the BZ (assuming the wrapping around all EPs follows the same circulation, such that all individual contributions to $W$ have the same sign). In particular, if a single EP is encircled once then one has $|W|=1/2$ [51, 52, 53] and, analogously, the eigenstates of both bands (assuming a two-band model) are also exchanged after traversing the BZ [51, 54]. This non-Hermitian $|W|=1/2$ case has been associated with the presence of a single edge state under open boundary conditions (OBC), and each $1/2$ increment in $|W|$ has been further associated with the appearance of an extra edge state [55]. Concerning the effective model for the single-impurity case studied in this section, it corresponds, to the best of our knowledge, to the first example of a Hermitian model with a half-integer winding number.
@@
Figure 7: (a) Lieb lattice with a mid-length line of onsite impurity potentials $\epsilon$ at every other site (red dots) along the $y$-direction. All hopping terms are given by $t$. (b) Diamond chain obtained from Fourier transforming the lattice in (a) along the periodic $y$-direction, using the indicated supercell. The $y$-momentum $k_{y}$ becomes the equivalent of a flux per plaquette. The dashed hoppings pick up a positive $k_{y}$ phase in the clockwise direction within each plaquette.
@@
From the examples mentioned above, the same pattern seems to emerge both for Hermitian and non-Hermitian models, namely that a half-integer bulk topological invariant is connected to the presence of a single edge state under OBC. The question now becomes: given that the same half-integer invariants (winding number and Zak phase) are observed for our single-impurity case, is there a physical signature of this non-trivial behavior? Recall that the “bulk” bands in Fig. 2(c) are defined over the reduced flux $\phi$, and not over a momentum variable $k$. Therefore, the concept of OBC does not directly apply here  [^1] Note that the bulk-boundary correspondence relates quantized bulk invariants defined for the momentum-space with the presence of edge states under OBC in the complementary real-space, since position and momentum are conjugate variables. The conjugate variable of the (reduced) magnetic flux is the (reduced) electric charge, a central relation in the study of quantum electrical circuits [78]. However, and for our purposes, it is not clear what the meaning of OBC in a “charge-space” could be, if any, and what would correspond to an edge state in this quantized charge-space.. To circumvent this limitation, we adopt a strategy based on an exact mapping as follows. First, we consider the modified Lieb lattice depicted in Fig. 7(a), with a line of impurity sites at every other site along the $y$-direction (red sites), positioned at the center of the $x$ coordinate. Then, assuming PBC (OBC) in the $y$-direction ($x$-direction), we Fourier transform the lattice along the $y$-direction (with lattice constant set to unity) using the supercell represented in Fig. 7(a), obtaining the open diamond chain in Fig. 7(b). As a result, this chain becomes the same as the single-impurity case of Fig. 1(a), with $k_{y}$ taking the place of $\phi$, and up to an inconsequential gauge transformation that accumulates all the $k_{y}$ “flux” in a single hopping term within each plaquette. Therefore, for a small value of $\epsilon$ we expect the effective low-energy subspace to behave as in Fig. 2(c), with $\phi\to k_{y}$. The same half integer invariants $\gamma$ and $W$ in (50) and (52), respectively, are obtained when integrating now over the Brillouin zone in $k_{y}$.
@@
Figure 8: (a) Energy spectrum, with $t$ as the energy unit, of the impurity decorated open Lieb lattice, with $N_{x}=41$ ($N_{y}=200$) unit cells in the $x$ ($y$) direction and impurities $\epsilon=0.1$ placed along unit cell 21 in $x$ according to the pattern in Fig. 7(a). (b) Zoomed energy spectrum of the region delimited by the orange dashed box in (a). (c)-(e) Depiction of the amplitude distribution of the eigenstate with index (c) $n=16201$, (d) $n=16202$, and (e) $n=16396$, with energies given by the corresponding encircled states in (b).
@@
In order to test if this translates into the appearance of a gapped edge state in real-space, we impose OBC along the $y$-direction also, i.e., we consider the Lieb lattice in Fig. 7(a) with OBC for both directions. In Fig. 8(a), we plot the energy spectrum for a chain $N_{x}=41$ ($N_{y}=200$) unit cells in the $x$ ($y$) direction, with impurities $\epsilon=0.1$ placed along unit cell 21 in $x$ according to the pattern in Fig. 7(a). The overall features of the spectrum in Fig. 8(a) are the expected ones for a Lieb lattice, namely a gapless system of two continua of dispersive states above and below the localized states of the zero-energy flat band. However, a closer look at the states with energies around $0\leq E\lesssim\frac{\epsilon}{2}$ [notice how these limits correspond to the energies of the flat bands of the effective system in Fig. 2(c)], shown in the zoomed region of Fig. 8(b), reveals the presence of an energy gap. As expected, an isolated edge state is seen to appear within this gap, with the spatial profile depicted in Fig. 8(c).
@@
The results of this exact mapping provide another example where the half-integer topological bulk invariants, $\gamma=\pi W=\frac{\pi}{2}$, have a correspondence with the presence of a single gapped edge state under OBC. In turn, this provides indirect evidence that the topologically non-trivial nature of the single-impurity case studied in this section can be related to specific physical manifestations. We note that this bulk-boundary correspondence should also be observed upon switching the single impurity from the $B$ site of the midchain plaquette, as in Fig. 7(b), to its $C$ site. In terms of the Lieb lattice in Fig. 7(a), this corresponds to shifting the impurities along the impurity line, exchanging its red and green sites. As we detail in Appendix A, this shift leads to a physically distinct configuration that produces small changes to the low-energy spectrum in Fig. 8(b). However, as we show there, the validity of this mapping is not affected, as a single in-gap state is still present.
V Conclusions
We studied a diamond chain with a gapped FB in its energy spectrum upon introducing a finite magnetic flux per plaquette, whose states can be described as CLSs spanning a non-orthogonal basis subspace. After outlining the mathematical formalism required for the analytical treatment of such non-orthogonal bases, we applied it to the general solutions of the impurity states that are lifted from the FB by small local potentials placed at the midchain plaquette. For the two equal, two symmetric and single-impurity cases, the numerical results were shown to closely follow the analytical predictions, except for flux values very close to the critical gap closing point between flat and dispersive bands, corresponding to the expected breakdown region of the theory.
@@ V Conclusions
A topological characterization of the impurity levels was carried out, where the flux was interpreted as a synthetic momentum. While the equal and symmetric impurity cases were characterized as topologically trivial, the single-impurity case manifested a peculiar type of non-trivial topology, namely one defined by half-integer invariants such as the Zak phase and winding number. The latter corresponds, as far as we know, to the first of its kind in the context of Hermitian models. Given that the invariants are computed by considering the reduced magnetic flux as the “bulk” variable of integration, and the variation of the flux by a flux quantum as the “Brillouin zone”, the bulk-boundary correspondence cannot be directly checked. However, an indirect way of circumventing this limitation was formulated, based on an exact mapping between the diamond chain with a single impurity and a two-dimensional Lieb lattice with a central line of impurities. The mapped lattice displays the same low-energy effective impurity bulk bands, classified by the same half-integer invariants. For fully open boundaries, its energy spectrum reveals the presence of a single edge state within the gap of this low-energy subspace, providing an indirect evidence of the unconventional bulk-boundary correspondence associated with the topology of the single-impurity case.
Regarding the realization of the model studied here, shown in Fig. 1(a), we highlight two platforms. One is electrical circuits [59], whose flexibility and plasticity allows one to recreate most tight-binding models simply by coupling nodes with capacitors or inductors in the desired geometry. By incorporating negative impedance converters with current inversion (INICs) [60, 61] in phase-control units, it has been shown [62] that arbitrary phase factors can be associated to the couplings, enabling the generation of a synthetic flux at each plaquette, as required by the target model of Fig. 1(a). The impurity potentials at the central nodes can be simulated by adding extra capacitors, with small capacitance values, to the grounding schemes of these nodes. Preliminary numerical studies should be undertaken to determine the optimal parameters of the circuit since, on the one hand, the admittance elements of the impurity nodes should be small, compared to the admittance gap between flat and dispersive bands, and, on the other hand, they should be sufficiently higher than the tolerance of the other components, so as to not have its effects washed way by this inherent disorder. As such, very low tolerance components should be considered. After feeding the circuit with an alternated current with the eigenfrequency of the impurity states, standard voltage and two-point impedance measurements can be used to detect them.
@@ Acknowledgments
A.M.M. and R.G.D. developed their work within the scope of the Portuguese Institute for Nanostructures, Nanomodelling, and Nanofabrication (i3N) Projects No. UIDB/50025/2020, No. UIDP/50025/2020, and No. LA/P/0037/2020, financed by national funds through the Fundação para a Ciência e Tecnologia (FCT) and the Ministério da Educação e Ciência (MEC) of Portugal. A.M.M. acknowledges financial support from i3N through the work Contract No. CDL-CTTRI-46-SGRH/2022. D.V. and V.A. acknowledge financial support from the Spanish State Research Agency AEI (contract No. PID2020-118153GBI00/AEI/10.13039/501100011033) and Generalitat de Catalunya (Contract No. SGR2021-00138).
Appendix A Comments on the single-impurity case
@@
where $M_{y}$ is the $y$-mirror operator.
Figure 9: (a) Energy spectrum, with $t$ as the energy unit, of the impurity decorated open Lieb lattice, with $N_{x}=41$ ($N_{y}=200$) unit cells in the $x$ ($y$) direction and impurities $\epsilon=0.1$ placed along unit cell 21 in $x$ as in Fig. 7(a), but with every impurity shifted by one position along the impurity line (corresponding to exchanging the red and green sites along this line). (b) Zoomed energy spectrum of the region delimited by the orange dashed box in (a). (c)-(e) Depiction of the amplitude distribution of the eigenstate with index (c) $n=16201$, (d) $n=16202$, and (e) $n=16396$, with energies given by the corresponding encircled states in (b).
Under this quasi-symmetry, the flux is reversed and the impurities exchange their positions, such that it only becomes an exact symmetry for the equal impurities case $\epsilon_{{}_{B}}=\epsilon_{{}_{C}}$. On the other hand, under time-reversal, the system behaves as
@@
Through the mapping developed for the single-impurity case in Sec. IVC, we arrived at the substitution $\phi\to k_{y}$. The corresponding combined quasi-symmetry acting on the diamond chain depicted in Fig. 7(b)) becomes, from (56), and up to a gauge transformation that redistributes the $k_{y}$ “flux” equally between all hopping terms of each plaquette,
@@
The question we want to address here is the following: when changing the impurities from the $B$ sites along the impurity line, as in Fig. 7(a), to the $C$ sites (by exchanging the red and green sites along this line), do the relevant qualitative features of the low-energy spectrum remain the same? In other words, does one still observe an energy gap between two flat or nearly-flat bands with a single in-gap edge state, as in Fig. 8(b), such that the bulk-boundary correspondence is preserved, that is, is the system insensitive to the choice of impurity positioning? In Fig. 9, we replicate the energy spectrum of the same Lieb lattice with the spectrum of Fig. 8(a), apart from changing the position of the impurities to the $C$ sites along the impurity line, following the procedure explained above. As can be seen in the zoomed low-energy spectrum of Fig. 9(b), a single in-gap state is found, albeit much closer in energy to the top bulk band than in Fig. 8(b). The spatial profile of this state is depicted in Fig. 8(c), confirming its edge localized nature. The energy state immediately above this one, depicted in Fig. 9(d), is confirmed to be a bulk state of the impurity line. Even though this first bulk state of the top band is closer to the in-gap state than in Fig. 8(b), a weaker hybridization with the edge state than for the state in Fig. 8(d) is observed. Finally, the last state of the top band is shown in Fig. 9(e), closely replicating the one in Fig. 8(e). In summary, while small quantitative differences are observed between the two choices of impurity positioning, both share the same qualitative features, most notably the presence of a single in-gap state in the relevant low-energy sector of the spectrum.

### v2

@@ Results
This behavior is similar to that observed in chiral-symmetric non-Hermitian 1D models where, due to the fact that each exceptional point (EP) [51] carries a topological charge of ±$\pm 1/2$, the absolute value of the winding number, $|W|$, corresponds to half the number of times each EP is encircled as the momentum $k$ traverses the BZ (assuming the wrapping around all EPs follows the same circulation, such that all individual contributions to $W$ have the same sign). In particular, if a single EP is encircled once then one has $|W|=1/2$ [52, 53, 54] and, analogously, the eigenstates of both bands (assuming a two-band model) are also exchanged after traversing the BZ [52, 55]. This non-Hermitian $|W|=1/2$ case has been associated with the presence of a single edge state under open boundary conditions (OBC), and each $1/2$ increment in $|W|$ has been further associated with the appearance of an extra edge state [56]. Concerning the effective model for the single-impurity case studied in this section, it corresponds, to the best of our knowledge, to the first example of a Hermitian model with a half-integer winding number.
@@
Figure 7: Mapping between Lieb lattice and diamond chain. (a) Lieb lattice with a mid-length line of onsite impurity potentials $\epsilon$ at every other site (red dots) along the $y$-direction. All hopping terms are given by $t$. (b) Diamond chain obtained from Fourier transforming the lattice in (a) along the periodic $y$-direction, using the indicated supercell. The $y$-momentum $k_{y}$ becomes the equivalent of a flux per plaquette. The dashed hoppings pick up a positive $k_{y}$ phase in the clockwise direction within each plaquette.
@@
From the examples mentioned above, the same pattern seems to emerge both for Hermitian and non-Hermitian models, namely that a half-integer bulk topological invariant is connected to the presence of a single edge state under OBC. The question now becomes: given that the same half-integer invariants (winding number and Zak phase) are observed for our single-impurity case, is there a physical signature of this non-trivial behavior? Recall that the “bulk” bands in Fig. 2(c) are defined over the reduced flux $\phi$, and not over a momentum variable $k$. Therefore, the concept of OBC does not directly apply here. Note that the bulk-boundary correspondence relates quantized bulk invariants defined for the momentum-space with the presence of edge states under OBC in the complementary real-space, since position and momentum are conjugate variables. The conjugate variable of the (reduced) magnetic flux is the (reduced) electric charge, a central relation in the study of quantum electrical circuits [58]. However, and for our purposes, it is not clear what the meaning of OBC in a “charge-space” could be, if any, and what would correspond to an edge state in this quantized charge-space. To circumvent this limitation, we adopt a strategy based on an exact mapping as follows. First, we consider the modified Lieb lattice depicted in Fig. 7(a), with a line of impurity sites at every other site along the $y$-direction (red sites), positioned at the center of the $x$ coordinate. Then, assuming PBC (OBC) in the $y$-direction ($x$-direction), we Fourier transform the lattice along the $y$-direction (with lattice constant set to unity) using the supercell represented in Fig. 7(a), obtaining the open diamond chain in Fig. 7(b). As a result, this chain becomes the same as the single-impurity case of Fig. 1(a), with $k_{y}$ taking the place of $\phi$, and up to an inconsequential gauge transformation that accumulates all the $k_{y}$ “flux” in a single hopping term within each plaquette. Therefore, for a small value of $\epsilon$ we expect the effective low-energy subspace to behave as in Fig. 2(c), with $\phi\to k_{y}$. The same half integer invariants $\gamma$ and $W$ in (50) and (52), respectively, are obtained when integrating now over the Brillouin zone in $k_{y}$.
@@
Figure 8: Lieb lattice with central impurity line. (a) Energy spectrum, with $t$ as the energy unit, of the impurity decorated open Lieb lattice, with $N_{x}=41$ ($N_{y}=200$) unit cells in the $x$ ($y$) direction and impurities $\epsilon=0.1t$ placed along unit cell 21 in $x$ according to the pattern in Fig. 7(a). (b) Zoomed energy spectrum of the region delimited by the orange dashed box in (a). The inset shows the energy spectrum of the zoomed region for the same Lieb lattice, only with PBC in the $y$-direction. (c)-(e) Depiction of the amplitude distribution, normalized by the maximum value for each case, of the eigenstate with index (c) $n=16201$, (d) $n=16202$, and (e) $n=16396$, with energies given by the corresponding encircled states in (b).
@@
In order to test if this translates into the appearance of a gapped edge state in real-space, we impose OBC along the $y$-direction also, i.e., we consider the Lieb lattice in Fig. 7(a) with OBC for both directions. In Fig. 8(a), we plot the energy spectrum for a chain $N_{x}=41$ ($N_{y}=200$) unit cells in the $x$ ($y$) direction, with impurities $\epsilon=0.1t$ placed along unit cell 21 in $x$ according to the pattern in Fig. 7(a). The overall features of the spectrum in Fig. 8(a) are the expected ones for a Lieb lattice, namely a gapless system of two continua of dispersive states above and below the localized states of the zero-energy flat band. However, a closer look at the states with energies around $0\leq E\lesssim\frac{\epsilon}{2}$ [notice how these limits correspond to the energies of the flat bands of the effective system in Fig. 2(c)], shown in the zoomed region of Fig. 8(b), reveals the presence of an energy gap. As expected, an isolated edge state is seen to appear within this gap, with the spatial profile depicted in Fig. 8(c). Conversely, when PBC are imposed in the $y$-direction, the in-gap state disappears, as shown in the inset of Fig. 8(b), which further corroborates its topological origin.
@@
The results of this exact mapping provide another example where the half-integer topological bulk invariants, $\gamma=\pi W=\frac{\pi}{2}$, have a correspondence with the presence of a single gapped edge state under OBC. In turn, this provides indirect evidence that the topologically non-trivial nature of the single-impurity case studied in this section can be related to specific physical manifestations. We note that this bulk-boundary correspondence should also be observed upon switching the single impurity from the $B$ site of the midchain plaquette, as in Fig. 7(b), to its $C$ site. In terms of the Lieb lattice in Fig. 7(a), this corresponds to shifting the impurities along the impurity line, exchanging its red and green sites. As we detail in Supplementary Note 1, this shift leads to a physically distinct configuration that produces small changes to the low-energy spectrum in Fig. 8(b). However, as we show in Supp. Fig. 1, the validity of this mapping is not affected, as a single in-gap state is still present.
Discussion
We studied a diamond chain with a gapped FB in its energy spectrum upon introducing a finite magnetic flux per plaquette, whose states can be described as CLSs spanning a non-orthogonal basis subspace. After outlining the mathematical formalism required for the analytical treatment of such non-orthogonal bases, we applied it to the general solutions of the impurity states that are lifted from the FB by small local potentials placed at the midchain plaquette. For the two equal, two opposite and single-impurity cases, the numerical results were shown to closely follow the analytical predictions, except for flux values very close to the critical gap closing point between flat and dispersive bands, corresponding to the expected breakdown region of the theory.
@@ Discussion
A topological characterization of the impurity levels was carried out, where the flux was interpreted as a synthetic momentum. While the equal and opposite impurity cases were characterized as topologically trivial, the single-impurity case manifested a peculiar type of non-trivial topology, namely one defined by half-integer invariants such as the Zak phase and winding number. The latter corresponds, as far as we know, to the first of its kind in the context of Hermitian models. Given that the invariants are computed by considering the reduced magnetic flux as the “bulk” variable of integration, and the variation of the flux by a flux quantum as the “Brillouin zone”, the bulk-boundary correspondence cannot be directly checked. However, an indirect way of circumventing this limitation was formulated, based on an exact mapping between the diamond chain with a single impurity and a two-dimensional Lieb lattice with a central line of impurities. The mapped lattice displays the same low-energy effective impurity bulk bands, classified by the same half-integer invariants. For fully open boundaries, its energy spectrum reveals the presence of a single edge state within the gap of this low-energy subspace, providing an indirect evidence of the unconventional bulk-boundary correspondence associated with the topology of the single-impurity case.
Regarding the realization of the model studied here, shown in Fig. 1(a), we highlight two platforms. One is electrical circuits [60], whose flexibility and plasticity allows one to recreate most tight-binding models simply by coupling nodes with capacitors or inductors in the desired geometry [61]. By incorporating negative impedance converters with current inversion (INICs) [62, 63] in phase-control units, it has been shown [64] that arbitrary phase factors can be associated to the couplings, enabling the generation of a synthetic flux at each plaquette, as required by the target model of Fig. 1(a). The impurity potentials at the central nodes can be simulated by adding extra capacitors, with small capacitance values, to the grounding schemes of these nodes. Preliminary numerical studies should be undertaken to determine the optimal parameters of the circuit since, on the one hand, the admittance elements of the impurity nodes should be small, compared to the admittance gap between flat and dispersive bands, and, on the other hand, they should be sufficiently higher than the tolerance of the other components, so as to not have its effects washed way by this inherent disorder. As such, very low tolerance components should be considered. After feeding the circuit with an alternated current with the eigenfrequency of the impurity states, standard voltage and two-point impedance measurements can be used to detect them.
@@
Data Availability
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
@@ Acknowledgments
A.M.M. and R.G.D. developed their work within the scope of the Portuguese Institute for Nanostructures, Nanomodelling, and Nanofabrication (i3N) Projects No. UIDB/50025/2020, No. UIDP/50025/2020, and No. LA/P/0037/2020, financed by national funds through the Fundação para a Ciência e Tecnologia (FCT) and the Ministério da Educação e Ciência (MEC) of Portugal. A.M.M. acknowledges financial support from i3N through the work Contract No. CDL-CTTRI-46-SGRH/2022, and from FCT through the work Contract No. CDL-CTTRI-91-SGRH/2024. D.V. and V.A. acknowledge financial support from the Spanish State Research Agency AEI (contract No. PID2020-118153GBI00/AEI/10.13039/501100011033) and Generalitat de Catalunya (Contract No. SGR2021-00138).
Author contributions
A.M.M. and D.V. developed the concept, analyzed the results and compared them with theoretical predictions. D.V. performed the numerical simulations. R.G.D. and A.M.M. developed the theoretical framework. A.M.M. wrote a first draft of the manuscript. V.A. and R.G.D. supervised the project. All authors discussed the results, read and revised the manuscript.
Competing Interests
The authors declare no competing interests.
Supplementary Note 1: Comments on the single-impurity case
@@
where $M_{y}$ is the $y$-mirror operator. Under this quasi-symmetry, the flux is reversed and the impurities exchange their positions, such that it only becomes an exact symmetry for the equal impurities case $\epsilon_{{}_{B}}=\epsilon_{{}_{C}}$. On the other hand, under time-reversal, the system behaves as
@@
Through the mapping developed for the single-impurity case in Sec. Impurity statesC, we arrived at the substitution $\phi\to k_{y}$. The corresponding combined quasi-symmetry acting on the diamond chain depicted in Fig. 7(b)) becomes, from (56), and up to a gauge transformation that redistributes the $k_{y}$ “flux” equally between all hopping terms of each plaquette,
@@
Figure 9: Lieb lattice with modified central impurity line. (a) Energy spectrum, with $t$ as the energy unit, of the impurity decorated open Lieb lattice, with $N_{x}=41$ ($N_{y}=200$) unit cells in the $x$ ($y$) direction and impurities $\epsilon=0.1t$ placed along unit cell 21 in $x$ as in Fig. 7(a), but with every impurity shifted by one position along the impurity line (corresponding to exchanging the red and green sites along this line). (b) Zoomed energy spectrum of the region delimited by the orange dashed box in (a). (c)-(e) Depiction of the amplitude distribution, normalized by the maximum value for each case, of the eigenstate with index (c) $n=16201$, (d) $n=16202$, and (e) $n=16396$, with energies given by the corresponding encircled states in (b).
The question we want to address here is the following: when changing the impurities from the $B$ sites along the impurity line, as in Fig. 7(a), to the $C$ sites (by exchanging the red and green sites along this line), do the relevant qualitative features of the low-energy spectrum remain the same? In other words, does one still observe an energy gap between two flat or nearly-flat bands with a single in-gap edge state, as in Fig. 8(b), such that the bulk-boundary correspondence is preserved, that is, is the system insensitive to the choice of impurity positioning? In Supp. Fig. 9, we replicate the energy spectrum of the same Lieb lattice with the spectrum of Fig. 8(a), apart from changing the position of the impurities to the $C$ sites along the impurity line, following the procedure explained above. As can be seen in the zoomed low-energy spectrum of Supp. Fig. 9(b), a single in-gap state is found, albeit much closer in energy to the top bulk band than in Fig. 8(b). The spatial profile of this state is depicted in Supp. Fig. 9(c), confirming its edge localized nature. The energy state immediately above this one, depicted in Supp. Fig. 9(d), is confirmed to be a bulk state of the impurity line. Even though this first bulk state of the top band is closer to the in-gap state than in Fig. 8(b), a weaker hybridization with the edge state than for the state in Fig. 8(d) is observed. Finally, the last state of the top band is shown in Supp. Fig. 9(e), closely replicating the one in Fig. 8(e). In summary, while small quantitative differences are observed between the two choices of impurity positioning, both share the same qualitative features, most notably the presence of a single in-gap state in the relevant low-energy sector of the spectrum.